The Wet Corp. has an investment project that will reduce expenses by $25,000 per year for 3 years. The project's cost is $30,000. If the asset is part of the 3-year MACRS category (33.33% first year depreciation) and the company's tax rate is 38%, what is the cash flow from the project in year 1? (Do not round intermediate calculations. Round your answer to the nearest dollar amount.)

$20,760

$20,080

$19,300

$18,750
Answer #1

Depreciation for year 1=30,000*33.33%=9999

Hence cash flow for project in year 1=Reduction in expenses*(1-tax rate)+Tax saving on Depreciation for year 1

=25,000*(1-0.38)+(0.38*9999)

=$19,300(Approx).